Hi need help with this differential equation problem. Thank you!

An integrating factor for the linear differential equation xy – y = x² is a) 0 b) 1 c) x d) 1/x e) e-x
Answer #1

Doubt or problrm in this then comment below.. i will explain you..

.

please thumbs up for this solution..thanks..

.

integrating factor = e^(-ln(x)) = 1 /x

answer = option d ) = 1/x ....